OpenStudy (anonymous):

Let an integer be chosen at random from the integers 1 to 30 inclusive. Find the probability that the integer chosen is divisible by 3. 9/30 11/30 1/3

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Um, what numbers are divisible by 3 between (and including) 1 and 30? Just find the first 10 multiples of 3.

OpenStudy (igreen):

3 * 1 = ? 3 * 2 = ? 3 * 3 = ? 3 * 4 = ? 3 * 5 = ? 3 * 6 = ? 3 * 7 = ? 3 * 8 = ? 3 * 9 = ? 3 * 10 = ?

OpenStudy (igreen):

Well, we don't even have to find them..just count how many there are.

OpenStudy (anonymous):

Sorry net is moving slow, or maybe it's the site. :/ 3,6,9,12,15,18,21,24,27,30,33,36

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Okay, count the one's between 1 and 30, how many are there? Yeah, it's Open Study's site.

OpenStudy (anonymous):

11, correct?

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

No..check again.

OpenStudy (anonymous):

9? I thought I was supposed to include 1 and 30.

OpenStudy (igreen):

No, 9 is wrong too..1 is not a multiple of 3.

OpenStudy (igreen):

1 is not divisible by 3*

OpenStudy (anonymous):

3 * 1 = 3 3 * 2 = 6 3 * 3 = 9 3 * 4 = 12 3 * 5 = 15 3 * 6 = 18 3 * 7 = 21 3 * 8 = 24 3 * 9 = 27 3 * 10 = 30 @igreen

OpenStudy (igreen):

Yep, and how many is that?

OpenStudy (anonymous):

You gave me 10 problems. Or are you talking about the 30?

OpenStudy (igreen):

Yes, 10 problems..

OpenStudy (igreen):

Since there are 10 numbers that are divisible by 3 between (and including) 1 and 30, the probability of picking a number divisible by 3 is 10/30. You can simplify 10/30, and what is that?

OpenStudy (anonymous):

1/3

OpenStudy (igreen):

Yep, so that's your answer.

OpenStudy (anonymous):

Thanks Green!